OpenStudy (anonymous):

fFind the exact circumference of a circle with the given radius. x/2

OpenStudy (anonymous):

\[ \pi D= \pi x \]

OpenStudy (anonymous):

If the radius is x/2, \[C = 2r*\pi\]\[C = 2(\frac x2)*\pi\]\[C = x*\pi\]

OpenStudy (anonymous):

thx guys appreciate it